### v1

@@ 2.2.3 Analysis in the spectral and nuclear norms
Consider the setting of Theorem 2.1 with an oversampling parameter $p\geq 2$, it holds that
$\mathbb{E}[\|\bm{A}-\widehat{\bm{A}}\|_{2}]$ $\leq\left(1+\frac{3k}{p-1}\beta_{k}^{(2)}+3\delta_{k}^{(2)}\right)\|\bm{\Sigma}_{2}\|_{2}+\frac{3e^{2}(k+p)}{p^{2}-1}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*},$ (15a)
$\mathbb{E}[\|\bm{A}-\widehat{\bm{A}}\|_{*}]$ $\leq\left(1+\frac{k}{p-1}\beta_{k}^{(*)}+\delta_{k}^{(*)}\right)\|\bm{\Sigma}_{2}\|_{*}.$ (15b)
Proof 2.5.
We start by proving Eq. 15a and use the structural bound (7) for the Nyström method to obtain
@@
Similarly to the proof of Section 2.2.1, we use the conditional distribution of $\bm{\Omega}_{2}|\bm{\Omega}_{1}$ and the triangle inequality for the $L_{2}$ norm twice, to get
@@
$\leq\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{2}+\big{(}\mathbb{E}_{\bm{\Omega}_{1}}\big{[}\mathbb{E}_{\bm{\Psi}}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}\mid\bm{\Omega}_{1}\big{]}\big{]}\big{)}^{1/2},$
where $\bm{\Psi}$ is an $(n-k)\times(k+p)$ standard Gaussian matrix. Then Eq. 32c leads to
@@
After taking expectation with respect to $\bm{\Omega}_{1}$ and using the triangle inequality for the $L_{2}$ norm, we have
@@
$\leq\sqrt{\frac{k}{p-1}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2}}+\sqrt{\delta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2}}+\sqrt{\frac{e^{2}(k+p)}{p^{2}-1}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*}}.$
We conclude the proof of Eq. 15a using the inequality $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$.
@@
$\leq\|\bm{\Sigma}_{2}\|_{*}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}^{1/2}_{22.1}\|_{\textup{F}}^{2}\frac{\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})}{p-1}$
$\leq\left(1+\frac{k}{p-1}\beta_{k}^{(*)}+\delta_{k}^{(*)}\right)\|\bm{\Sigma}_{2}\|_{*}.$
@@
Using the notation of Theorem 2.1 with an oversampling parameter $p\geq 4$, and $u,t\geq 1$, it holds that
$\|\bm{A}-\widehat{\bm{A}}\|_{2}$ $\leq\left(1+4\delta_{k}^{(2)}+4(d_{1}+d_{2}u^{2})t^{2}\beta_{k}^{(2)}\right)\|\bm{\Sigma}_{2}\|_{2}+4d_{2}t^{2}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*},$ (16a)
$\|\bm{A}-\widehat{\bm{A}}\|_{*}$ $\leq\left(1+2\delta_{k}^{(*)}+d_{1}t^{2}\beta_{k}^{(*)}\right)\|\bm{\Sigma}_{2}\|_{*}+2d_{2}t^{2}u^{2}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2},$ (16b)
where each inequality holds with probability $\geq 1-2t^{-p}-e^{u^{2}/2}$.
Proof 2.7.
We begin by deriving the tailbound Eq. 16a in the spectral norm. Similar to Section 2.2.2, we process the second term of the structural bound Eq. 7 with the concentration inequality of Lemma 2.2 in the spectral norm:
$\mathbb{P}\left\{\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}\geq\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}\mid\bm{\Omega}_{1}\big{]}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}\|\bm{\Omega}_{1}^{\dagger}\|_{2}u\mid\bm{\Omega}_{1}\right\}\leq e^{-u^{2}/2}.$ (17)
Using Lemma A.3, it holds that
$\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}\mid\bm{\Omega}_{1}\big{]}$
$\leq\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{2}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}\|\bm{\Omega}_{1}^{\dagger}\|_{2}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}.$
@@
where $\mathbb{P}(E_{t}^{c})\leq 2t^{-p}$ by Lemma A.7. We conclude the proof of Eq. 16a by conditioning Eq. 17 on $E_{t}$ and using the inequality $\|\bm{x}\|_{1}\leq\sqrt{d}\|\bm{x}\|_{2}$ for $\bm{x}\in\mathbb{R}^{d}$ to obtain:
@@
with probability $\geq 1-2t^{-p}-e^{-u^{2}/2}$.
The proof of the tail bound for the nuclear norm Eq. 16b follows from a similar argument.
@@
$\mathbb{E}\|\bm{B}-\bm{Q}\bm{Q}^{*}\bm{B}\|_{\textup{F}}^{2}\leq\left(1+\frac{k}{p-1}\tilde{\beta}_{k}+\tilde{\delta}_{k}\right)\|\bm{S}_{2}\|_{\textup{F}}^{2},$
@@
where $\tilde{\beta}_{k}=\operatorname{tr}(\bm{S}_{2}\widetilde{\bm{K}}_{22.1}\bm{S}_{2})\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}/\|\bm{S}_{2}\|_{\textup{F}}^{2}$ and $\tilde{\delta}_{k}=\operatorname{tr}(\bm{S}_{2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{S}_{2})\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}/\|\bm{S}_{2}\|_{\textup{F}}^{2}$. This bound coincides with the standard randomized SVD bound [20, Thm. 10.5] when $\bm{K}=\bm{I}$, unlike the bound proved in [7, Thm. 2].

### v2

@@ 2.3.1 Expectation bounds
Let $\bm{A}$ be an $n\times n$ SPSD matrix, $2\leq k\leq\operatorname{rank}(\bm{A})$ be a target rank, and $p\geq 2$ an oversampling parameter. Let $\bm{\Omega}$ be a random sketch matrix with the columns i.i.d. $\mathcal{N}(\bm{0},\bm{K})$, where the covariance matrix $\bm{K}$ is such that the matrix $\widetilde{\bm{K}}_{11}$ defined in (3) is invertible. Then, the Nyström approximation $\widehat{\bm{A}}$ satisfies
(6a) $\mathbb{E}[\|\bm{A}-\widehat{\bm{A}}\|_{2}]$ $\leq\left(1+\frac{3k}{p-1}\beta_{k}^{(2)}+3\delta_{k}^{(2)}\right)\|\bm{\Sigma}_{2}\|_{2}+\frac{3e^{2}(k+p)}{p^{2}-1}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*},$
(6b) $\mathbb{E}[\|\bm{A}-\widehat{\bm{A}}\|_{*}]$ $\leq\left(1+\frac{k}{p-1}\beta_{k}^{(*)}+\delta_{k}^{(*)}\right)\|\bm{\Sigma}_{2}\|_{*}.$
Proof 2.2.
To prove the inequalities Eq. 6a and Eq. 6b, we use the $L_{q}$ norm of a random variable $Z$ defined as $\mathbb{E}^{q}(Z)=\mathbb{E}[|Z|^{q}]^{1/q}$. After taking expectation with respect to $\bm{\Omega}$ and applying Hölder’s inequality, the second term in the bound (5) for the operator norm satisfies
@@
To proceed from here, we recall that Section 2.1 provides the conditional distribution of $\bm{\Omega}_{2}|\bm{\Omega}_{1}$ as $\bm{\Omega}_{2}|\bm{\Omega}_{1}\sim\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}+\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}$, where $\bm{\Psi}$ is a standard Gaussian matrix. We use the conditional distribution of $\bm{\Omega}_{2}|\bm{\Omega}_{1}$, the triangle inequality for the $L_{2}$ norm twice, and $\bm{\Omega}_{1}\bm{\Omega}_{1}^{\dagger}=\bm{I}_{k}$ (which holds with probability $1$), to get
@@
$\leq\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{2}+\big{(}\mathbb{E}_{\bm{\Omega}_{1}}\big{[}\mathbb{E}_{\bm{\Psi}}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}\mid\bm{\Omega}_{1}\big{]}\big{]}\big{)}^{1/2}.$
Then Eq. 36c leads to
@@
After taking expectation with respect to $\bm{\Omega}_{1}$ and using Hölder’s inequality we have
@@
$\leq\sqrt{\frac{k}{p-1}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2}}+\sqrt{\delta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2}}+\sqrt{\frac{e^{2}(k+p)}{p^{2}-1}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*}},$
where we used $\|\widetilde{\bm{K}}_{11}^{-1/2}\|_{\textup{F}}\leq\sqrt{k}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{1/2}$ and the definitions of $\beta_{k}^{(2)}$, $\beta_{k}^{(*)}$, and $\delta_{k}^{(2)}$. We conclude the proof of Eq. 6a using the inequality $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$.
@@
${\color[rgb]{0,0,0}=}\|\bm{\Sigma}_{2}\|_{*}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}^{1/2}_{22.1}\|_{\textup{F}}^{2}\frac{\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})}{p-1}$
$\leq\left(1+\frac{k}{p-1}\beta_{k}^{(*)}+\delta_{k}^{(*)}\right)\|\bm{\Sigma}_{2}\|_{*},$
where we used (36a) for the equality and $\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})\leq k\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}$ for the inequality.
2.3.2 Tail bounds
We now proceed with the deriving tail bounds for $\|\bm{A}-\widehat{\bm{A}}\|$ in the operator and nuclear norms. We begin with the proof of a concentration inequality on norms of shifted and rescaled Gaussian matrices.
Lemma 2.3.
holds for every $u\geq 1$.
Proof 2.4.
Given $h(\bm{X}):=\|\bm{B}+\bm{C}\bm{X}\bm{D}\|_{(s)}$, we have
where we used that the Frobenius norm is larger than any Schatten-$s$ norm for $s\geq 2$. In other words, $h$ is Lipschitz continuous with Lipschitz constant $\|\bm{C}\|_{2}\|\bm{D}\|_{2}$. This allows us to apply a concentration result for functions of Gaussian matrices [21, Prop. 10.3], which yields the statement of the lemma.
With Lemma 2.3 at hand, we are ready to state and prove the tail bound.
@@ 2.3.2 Tail bounds
Using the notation of Theorem 2.8 with an oversampling parameter $p\geq 4$, define
(7) $d_{1}=\frac{3k}{p+1},\quad d_{2}=e^{2}\frac{k+p}{(p+1)^{2}}.$
Then for any $u,t\geq 1$, each of the two inequalities
(8a) $\|\bm{A}-\widehat{\bm{A}}\|_{2}$ $\leq\left(1+4\delta_{k}^{(2)}+4(d_{1}+d_{2}u^{2})t^{2}\beta_{k}^{(2)}\right)\|\bm{\Sigma}_{2}\|_{2}+4d_{2}t^{2}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*},$
(8b) $\|\bm{A}-\widehat{\bm{A}}\|_{*}$ $\leq\left(1+2\delta_{k}^{(*)}+d_{1}t^{2}\beta_{k}^{(*)}\right)\|\bm{\Sigma}_{2}\|_{*}+2d_{2}t^{2}u^{2}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2},$
holds with probability at least $1-2t^{-p}-e^{u^{2}/2}$.
Proof 2.6.
We begin by deriving the tail bound Eq. 8a in the spectral norm. To obtain Eq. 8a, it suffices to derive a tail bound for the term $\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(2)}^{2}$ in the structural bound (5). Using that $\bm{\Omega}_{2}|\bm{\Omega}_{1}\sim\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}+\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}$ with a standard Gaussian matrix $\bm{\Psi}$, Lemma 2.3 yields the following tail bound for $\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}$ conditioned on $\bm{\Omega}_{1}$:
(9) $\mathbb{P}\left\{\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}\geq\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}\mid\bm{\Omega}_{1}\big{]}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}\|\bm{\Omega}_{1}^{\dagger}\|_{2}u\mid\bm{\Omega}_{1}\right\}\leq e^{-u^{2}/2}.$
Using Lemma A.1 and the triangle inequality, it holds that
$\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}\mid\bm{\Omega}_{1}\big{]}\leq\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{2}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}\|\bm{\Omega}_{1}^{\dagger}\|_{2}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}.$
@@
where $d_{1}$ and $d_{2}$ are defined in (7) and $\mathbb{P}(E_{t}^{c})\leq 2t^{-p}$ by Lemma A.5. Then, conditioned on $E_{t}$, by Eq. 9 and the inequality $\|\bm{x}\|_{1}\leq\sqrt{d}\|\bm{x}\|_{2}$ for $\bm{x}\in\mathbb{R}^{d}$ we have:
@@
with probability $\geq 1-e^{-u^{2}/2}$. Similarly to the proof of [21, Thm. 10.8], we remove the conditioning by a union bound using $\mathbb{P}(E_{t}^{c})\leq 2t^{-p}$ and conclude the proof of (8a). The proof of the tail bound for the nuclear norm Eq. 8b follows from a similar argument.
@@
(10) $\mathbb{E}\|\bm{B}-\bm{Q}\bm{Q}^{*}\bm{B}\|_{\textup{F}}^{2}\leq\left(1+\frac{k}{p-1}\tilde{\beta}_{k}+\tilde{\delta}_{k}\right)\|\bm{S}_{2}\|_{\textup{F}}^{2},$
where $\tilde{\beta}_{k}=\operatorname{tr}(\bm{S}_{2}\widetilde{\bm{K}}_{22.1}\bm{S}_{2})\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}/\|\bm{S}_{2}\|_{\textup{F}}^{2}$ and $\tilde{\delta}_{k}=\operatorname{tr}(\bm{S}_{2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-2}\widetilde{\bm{K}}_{21}^{*}\bm{S}_{2})/\|\bm{S}_{2}\|_{\textup{F}}^{2}$. In comparison, by following the analysis of the prior work in [7] one can show that
(11) ${\color[rgb]{0,0,0}\mathbb{E}\|\bm{B}-\bm{Q}\bm{Q}^{*}\bm{B}\|_{\textup{F}}^{2}\leq\left(1+\frac{k(k+p)}{p-1}\tilde{\gamma}_{k}\right)\|\bm{S}_{2}\|_{\textup{F}}^{2},}$
@@
where $\tilde{\gamma}_{k}=\operatorname{tr}(\bm{S}_{2}\widetilde{\bm{K}}_{22}\bm{S}_{2})\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}/\|\bm{S}_{2}\|_{\textup{F}}^{2}.$ Since $\tilde{\beta}_{k},\tilde{\delta}_{k}\leq\tilde{\gamma}_{k}$, the inequality (10) improves on (11). Furthermore, (10) coincides with the standard randomized SVD bound [21, Thm. 10.5] when $\bm{K}=\bm{I}$, unlike (11).
2.4 Analysis in the Frobenius norm
We now proceed with the error bounds for $\|\bm{A}-\widehat{\bm{A}}\|_{\textup{F}}$. These are technical, but important to derive error bounds in Hilbert-Schmidt norm in the continuous setting in Section 3. The Hilbert-Schmidt norm controls the variance of the stochastic trace estimator of integral operators, see [47], and is equal to the $L^{2}$-norm error of the kernel of the integral operators. The key ideas of the analysis will be similar to the analysis in Section 2.3, but requires bounding $\|(\bm{\Sigma}_{2}^{1/2}\bm{\bm{\Omega}}_{2}\bm{\Omega}_{1}^{\dagger})^{*}\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}=\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}$, which is more complicated.
2.4.1 Expectation bound
We begin with stating and proving the expectation bound.
Theorem 2.8 (Frobenius norm).
Consider the setting of Theorem 2.1 with an oversampling parameter $p\geq 4$. Define the following constants
(12) $c_{1}=k\frac{(p-1)(k+1)+2}{p(p-1)(p-3)},\quad c_{2}=k\frac{k+p-1}{p(p-1)(p-3)},$
Then, the Nyström approximation $\widehat{\bm{A}}$ satisfies
(13) $\mathbb{E}[\|\bm{A}-\widehat{\bm{A}}\|_{\textup{F}}]\leq\left(1+2\delta_{k}^{(\textup{F})}+2\sqrt{c_{1}}\beta_{k}^{(\textup{F})}\right)\|\bm{\Sigma}_{2}\|_{\textup{F}}+2\sqrt{c_{2}}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*}.$
Proof 2.9.
To prove inequality Eq. 13 of Theorem 2.8, we once again use the $L_{q}$ norm of a random variable $Z$ defined as $\mathbb{E}^{q}(Z)=\mathbb{E}[|Z|^{q}]^{1/q}$. After taking expectation with respect to $\bm{\Omega}$ and applying Hölder’s inequality, the second term in the bound (5) for the Frobenius norm satisfies
To proceed from here, we, once again, recall that Section 2.1 provides the conditional distribution of $\bm{\Omega}_{2}|\bm{\Omega}_{1}$ as $\bm{\Omega}_{2}|\bm{\Omega}_{1}\sim\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}+\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}$, where $\bm{\Psi}$ is a standard Gaussian matrix. Using the triangle inequality for the $L_{4}$ norm and $\bm{\Omega}_{1}\bm{\Omega}_{1}^{\dagger}=\bm{I}_{k}$ (which holds with probability $1$), we obtain that
(14) $\mathbb{E}^{4}[\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}]$ $=\mathbb{E}_{\bm{\Omega}_{1},\bm{\Psi}}^{4}[\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}\bm{\Omega}_{1}^{\dagger}+\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{(4)}]$
To bound the second term, we first apply Eq. 36b,
and then take expectation with respect to $\bm{\Omega}_{1}$ using Lemma A.1 and inequalities in Lemma A.3:
${\color[rgb]{0,0,0}\mathbb{E}_{\bm{\Omega}_{1}}\big{[}\mathbb{E}_{\bm{\Psi}}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}\mid\bm{\Omega}_{1}\big{]}\big{]}\leq\left(c_{1}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{4}+c_{2}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}^{4}\right)\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{2},}$
where $c_{1}$ and $c_{2}$ are defined in (12). Inserting the inequality above into (14) and using $\|\widetilde{\bm{K}}_{11}^{-1/2}\|_{2}=\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{1/2}$ gives
$\quad\qquad\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{(4)}+\big{(}c_{1}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{4}+c_{2}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}^{4}\big{)}^{1/4}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{1/2},$
where we used $(a+b)^{2}\leq 2a^{2}+2b^{2}$ and the subadditivity of the square-root. This concludes the proof of Eq. 13.
2.4.2 Tail bound
We now proceed with stating and proving the tail bound for the Frobenius norm error.
Theorem 2.10 (Frobenius norm).
Consider the setting of Theorem 2.1 with an oversampling parameter $p\geq 4$. Define $d_{3}=\sqrt{k}d_{2}$, where $d_{2}$ is defined in (7) and let $u,t\geq 1$. Then,
(15) $\|\bm{A}-\widehat{\bm{A}}\|_{\textup{F}}$ $\leq\|\bm{\Sigma}_{2}\|_{\textup{F}}+4\left(\delta_{k}^{(\textup{F})}+t^{2}(d_{1}+d_{3})\beta_{k}^{(\textup{F})}\right)\|\bm{\Sigma}_{2}\|_{\textup{F}}+4t^{2}d_{3}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*}$
$\quad+2t^{2}u^{2}d_{2}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2}$
holds with probability at least $1-3t^{-p}-e^{-u^{2}/2}$.
Proof 2.11.
To obtain Eq. 15, by (5) it suffices to derive a tail bound for $\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}$. Using that $\bm{\Omega}_{2}|\bm{\Omega}_{1}\sim\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}+\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}$ with a standard Gaussian matrix $\bm{\Psi}$, Lemma 2.3 yields the following tail bound for $\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}$ conditioned on $\bm{\Omega}_{1}$:
(16) $\mathbb{P}\left\{\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\geq\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\mid\bm{\Omega}_{1}\big{]}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}\|\bm{\Omega}_{1}^{\dagger}\|_{2}u\mid\bm{\Omega}_{1}\right\}\leq e^{-u^{2}/2}.$
For a deterministic matrix $\bm{X}$ and random matrix $\bm{Y}$, combining Jensen’s inequality and the triangle inequality yields $\mathbb{E}\|\bm{X}+\bm{Y}\|_{(4)}\leq\mathbb{E}^{4}\left[\|\bm{X}+\bm{Y}\|_{(4)}\right]\leq\|\bm{X}\|_{(4)}+\mathbb{E}^{4}\left[\|\bm{Y}\|_{(4)}\right]$. Hence, by Eq. 36b we get
${\color[rgb]{0,0,0}=}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{(4)}+\left(\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{4}\big{(}\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}+\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}\big{)}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}^{4}\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}\right)^{1/4}.$
(17) $\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\mid\bm{\Omega}_{1}\big{]}^{2}$ $\leq 2\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{(4)}^{2}+2\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{2}(\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}+\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{2})$
with the constants $d_{1},d_{2},$ and $d_{3}$ defined in (7) and in the statement of Theorem 2.10. Lemma A.5 together with the inequalities $\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})\leq k\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}$ and $\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}\leq\sqrt{k}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}$ implies the following bound for $\mathbb{P}(E_{t}^{c})$:
(18) $\mathbb{P}(E_{t}^{c})\leq 2t^{-(p+1)}+t^{-p}\leq 3t^{-p}.$
Under the event $E_{t}$, Eq. 17 becomes
(19) $\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\mid E_{t}\big{]}^{2}\leq 2\left(\delta_{k}^{(\textup{F})}+t^{2}(d_{1}+d_{3})\beta_{k}^{(\textup{F})}\right)\|\bm{\Sigma}_{2}\|_{\textup{F}}+2t^{2}d_{3}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*}.$
Conditioning Eq. 16 on $E_{t}$ and combining it with Eq. 19 yields
with probability $\geq 1-e^{-u^{2}/2}$. Once again, we remove the conditioning by a union bound using Eq. 18 and conclude the proof of Theorem 2.8.